OpenStudy (anonymous):

Divide. (-6m^7+20m^6) + 2m^4 A. -3m^7+10m^6 B. -6m^7+20m^2 C.-3m^3+10m^2 D. -3m^3+20m^6

OpenStudy (anonymous):

C.-3m^3+10m^2

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

ok im about to post my new question right here ok

OpenStudy (anonymous):

Simplify the expression \[\sqrt{6}+2\sqrt{6}\]

OpenStudy (anonymous):

=\[3\sqrt{6}\]

OpenStudy (anonymous):

ok im bout to start a new question and the thing will say posted below